# Capsule Network Projectors are Equivariant and Invariant Learners

### v2

@@ 4.1 Training Protocol
To directly compare with prior works employing the 3DIEBench dataset, we follow an identical training protocol, as defined in [14]. All methods employ a ResNet-18 encoder network ($f_{\theta}$), for the projection head ($h_{\phi}$) we compare a variety of hyperparameterisations, which we later describe in the following sections. For primary bench-marking we train our model for 2000 epochs using the Adam [19] optimiser with default settings, a fixed learning rate of 1e-3 and batch size of 1024. For ablations and sensitivity analyses we train for 500 epochs and employ a batch size of 512, with other settings remaining unchanged. We find in practice that 500 epochs presents a strong correlation of performance. By default the objective function weighting are as follows, $\lambda_{\text{inv}}=0.1,\;\lambda_{\text{equi}}=5,\;\lambda_{V}=10,\;\lambda_{C}=1$. Each self-supervised 2000 epoch pretraining run took approximately 22 hours using 3 Nvidia A100 80GB GPU’s, with 64 capsule models taking approximately 25 hours using 6 Nvidia A100 80GB GPU’s. All eval tasks are completed on a single Nvidia A100 80GB GPU and take approximately 6 hours for angle prediction, and 3 hours for classification.
Table 1: Evaluation of invariant properties on downstream classification task. We evaluate both the representations and the intermediate embeddings of the projection head when different numbers of capsules in the projection head is used. ‘-’ refers to non-compatible experiments.
@@
SIE [14] 20.1M 13.07G 512 512 82.94 82.08 80.32
CapsIE - 16 12.7M 3.49G 16 256 78.96 - -
CapsIE - 32 14.7M 4.57G 32 512 80.00 - -
CapsIE - 64 20.4M 8.69G 64 1024 80.26 - -
@@
CapsIE - 16 12.7M 3.49G 16 256 - 82.96 -
CapsIE - 32 14.7M 4.57G 32 512 - 83.49 -
CapsIE - 64 20.4M 8.69G 64 1024 - 83.64 -
@@ 4.2 Downstream Evaluation
To evaluate the quality of representations learnt under self-supervision, we use the standard benchmark approach of learning downstream task specific networks with frozen representations as input. In our case we evaluate the representations in three distinct tasks to evaluate both invariant and equivariant properties, we use a linear evaluation training protocol of the frozen representations. Further details of the evaluation protocol are given in the appendix B.2.
@@
Equivariant Evaluation. Evaluating equivariant properties is achieved through a rotation prediction task in which a three layer MLP is trained to predict the quaternions defining the rotation between two views of the same object. We train for 300 epochs using MSE loss. Similarly to rotation prediction, we evaluate the representation’s equivariant properties by regressing the the colour hue of an object view. We train a single linear layer for 50 epochs using MSE loss.
Representation Quality. The performance of CapsIE for both invariant and equivariant benchmark tasks is given in Tables 1 and 2 respectively. We evaluate both the representations produced by the ResNet-18 encoder and the intermediate embeddings of the capsule layer projection head given different values for the numbers of capsules. We observe that across all models that the invariant properties captured within the representations marginally suffer compared to the MLP projector of SIE. This observation is expected given the significantly reduced number of embeddings employed in the invariant criterion compared to SIE. However, the evaluation of equivariant properties captured by the representations demonstrates that the use of a capsule projector in place of a MLP can lead to vastly improved performance in rotation prediction ($\uparrow 0.05\;R^{2}$) advancing the prior state-of-the-art whist also improving on the supervised baseline by a significant margin.
Interestingly, we also observe that the colour prediction task achieves a performance close to that of the supervised setting even though our criteria do not directly optimise for such equivariant properties. This suggests that the CapsNet is responsible for implicitly capturing transformations of the input whilst having little to no impact on the tasks directly optimised for. In addition, to the improvement on SOTA our CapsIE model has the favourable consequence of being far more parameter and computationally efficient.
Table 2: Evaluation of equivariant properties on downstream rotation prediction (left) and colour prediction (right) tasks. We evaluate both the representations and the intermediate embeddings of the projection head when different numbers of capsules in the projection head is used.
@@
SIE [14] 0.73 0.23 0.73 0.07 0.05 0.02
CapsIE - 16 0.78 - - 0.97 - -
CapsIE - 32 0.75 - - 0.97 - -
CapsIE - 64 0.72 - - 0.97 - -
@@
CapsIE - 16 - - 0.78 - - 0.97
CapsIE - 32 - - 0.77 - - 0.97
CapsIE - 64 - - 0.78 - - 0.97
Intermediate Projector Embeddings. The role of the projector is primarily employed to decorrelate the embeddings on which the objective function operates on from the representations employed downstream. The premise is to avoid representations that are over-fit to the self-supervised objective [4]. However, it has been well studied that it can be beneficial to maintain a number of projector layers and instead utilise intermediate projector embeddings for downstream tasks. In our case, the preservation of capsule layers for downstream tasks ensures that the desirable equivariant and part-whole properties are maintained. Specifically, the equivariant information that has shown to be captured in the object pose [27].
[…]
@@ 4.3 Quantitative Evaluation of Equivariance
In order to quantitatively evaluate the equivariant performance of our method and capsule projector, we employ commonly used metrics including Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k). We utilise the same setup as described in [14], and discuss in further details the setup in appendix B.3. All the results evaluated by the aforementioned metrics are given in Table 3. Our CapsIE network outperforms EquiMod, Only Equivariance and SIE by a considerable margin across all metrics and for all dataset splits. We achieve strong perfromance on PRE, reporting 0.21 PRE on the validation set compared to 0.48 for EquiMod and Only Equivariance and 0.29 for SIE. The same large gains in equivariant performance are shown for MRR and H@1 and H@5. Note, a random H@1 results in a performance for 2% (0.02) demonstrating that our method lies well above random. Given that the same predictor is used for all reported methods, we can thus conclude how impactful the CapNet projector is at learning equivariant embeddings as opposed to linear projectors.
Table 3: Quantitative evaluation of the predictor when using a Capsule network projector, using PRE, MRR and H@k. The source dataset which embeddings are computed and the dataset used for retrieval are given in the format source-retrieval for PRE and source for MRR and H@k.
@@
CapsIE (ours) 0.17 0.21 0.20 0.60 0.47 0.50 0.36 0.71 0.58
5 Ablations & Sensitivity Analysis
@@ 5.1 Number of Capsules
Each capsule, in theory, should represent a unique concept, thus when the number of capsules is increased, logically so should the networks representation ability to capture an increasing number of semantic concepts. Observing the invariant performance during training (Figure 5) and downstream evaluation in Table 1, CapsIE gains slight improvement with the addition of more capsules. This demonstrates that our model has better utilised the additional representational power to improve performance. This pattern is also observed when evaluating equivariant properties (shown in Figure 5), yet is less pronounced. However, in Table 1 we also show that increasing the number of capsules of a supervised SR-Caps model trained in a standard supervised fashion is not an indicator of increased performance, aligning with prior capsule research [12]. This differentiation in behaviour provides an interesting direction for future research.
5.2 Implicit Viewpoint Equivariance
Table 4: Evaluation of SIE and CapsIE - 32 downstream rotation prediction on representations by either a rotation or colour hue equivariant objective.
Rotation Prediction ($R^{2}$)
Method All
SIE - Rotation 0.43
SIE - Colour 0.29 ($\downarrow$ 0.14)
CapsIE - Rotation 0.59
CapsIE - Colour 0.48 ($\downarrow$ 0.11)
To investigate whether capsule models are implicitly learning equivariant properties without any explicit enforcing criterion, we train our CapsIE model to predict colour hue rather than rotation via the predictor network. Given the observation in Table 2 that CapsIE is able to achieve near-perfect prediction without any optimising criterion, we hypothesise that capsules are implicitly capable of learning equivariant properties. The evaluation performance of representations on the rotation task under the aforementioned ablation pre-training settings in Table 5, demonstrates that CapsNets indeed learns more implicit equivariant properties than the benchmark SIE by a considerable margin. We do however observe that SIE is still able to capture some equivariant information without being explicitly trained for, which does not align with the findings of the setting where colour is evaluated [14]. Notably, this results is not a significant improvement over random representations in which the $R^{2}$ lies approximately at 0.25. This observation presents interesting study for future work, while the improvement of the CapsIE model over SIE empirically demonstrates the viewpoint equivariant assumptions of CapsNets.
Figure 4: Classification evaluation performance (top-1 %) during training for given varying number of capsules.
Figure 5: Rotation prediction evaluation performance ($R^{2}$) during training for given varying number of capsules.
@@ 6.1 Equivariant Self-Supervised Learning
Self-supervised learning has seen the majority of its success in the invariant setting by either contrastive [7], information maximisation [33, 3], or clustering methods [5, 1]. All families of approach rely on training a network to be invariant to transformations by increasing the similarity between embeddings of the same image under augmentation. The differing approaches emerge from alternative methods to avoid collapse, the phenomena where embeddings fall into a lower-dimensional subspace rather than the entire available embedding space resulting in a trivial solution [18]. Although these methods differ, they all produce similarly performing representations, hence we employ information maximisation methods as the basis of this work due to their computational efficiency.
Learning to be invariant to transformations is typically useful for semantic discrimination tasks, yet preserving information about the transformations can be highly beneficial. Some approaches have attempted to capture specific information regarding transformations by predicting the applied augmentation parameters [21], preserving the strength of augmentations [31] and introducing rotational transformations [8]. However, as stated in [14], these methods provide no guarantee that a mapping is learnt in the latent space that reflects the transformations in the input space. Hence, methods have been employed that address this limitation, [10, 26, 14] all employ predictor networks to predict the displacement representations in the latent space given one view representation and the transformation parameters. The latter, SIE [14], is the basis of our work, which further extends prior methods by splitting representation vectors into invariant and equivariant parts to better separate differing information.
@@ 6.2 Capsule Networks
Figure 6: Objects with similar semantic concepts (i.e. bus and car) have corresponding capsule activation/invariant embeddings. Top-3 capsule activations are highlighted.
CapsNets present an alternative architecture to CNNs, addressing their limitations by explicitly preserving hierarchical spatial relationships between features [29]. CapsNets replace scalar neurons with vector or matrix poses, representing specific concepts at different levels of a parse tree as the network goes deeper. The first layer (primary capsules) correspond to the most basic parts, while capsules in deeper layers represent more complex concepts made up of the simpler concepts as they get closer to the final layer where each capsule corresponds to a specific class (depicted in Figure 6).
@@
The key novelty in CapsNets is the routing mechanism, which determines the contributions of lower-level capsules to higher-level capsules. Numerous routing algorithms, both iterative and non-iterative, have been proposed to address the efficiency and effectiveness of this process [29, 23, 13, 28, 17, 9, 32, 11]. Among these, SRCaps [15] introduces a non-iterative routing mechanism. This method maintains all of the desirable properties of CapsNets, such as equivariance while largely mitigating the time cost of iterative methods, at the price of a small amount of performance. However, SRCaps faces the same limitations as other CapsNets where high resolution or high class datasets are beyond the network’s abilities when trained in a standard fashion. For a more detailed description of capsule routing mechanisms please check this review [27].
@@ 7 Conclusion
Our proposed method demonstrates how CapsNets can be employed in self-supervised learning to better learn equivariant properties without the need to explicitly split representation vectors and train separate projector networks. The resulting solution, CapsIE, achieves state-of-the-art performance in equivariant and invariant downstream benchmarks with a significant improvement of 0.05 $R^{2}$ on prior self-supervised rotation prediction tasks and 0.02 $R^{2}$ improvement over the supervised baseline. In addition, we observe competitive performance of CapsIE on equivariant tasks not explicitly trained for, further demonstrating the implicit equivariant properties of our capsule architecture under standard invariant optimisation criteria akin to those of VICReg [3], SimCLR [7], and MSN [1]. Our results contribute significantly to the avocation of CapsNets in self-supervised representation learning, introducing desirable properties with improved effectiveness over MLP projectors.
Limitations & Broader Impact. This work aims to learn higher quality and more applicable representations of images without human generated annotations, therefore such methods can lead to positive societal impacts the development of more accurate or informative models for a number of downstream tasks. However, as is the case with all vision systems, there is potential for exploitation and security concerns and one should take into consideration AI misuse when extending our method.
In addition, the use of CapsNets has shown little improvement when scaled, similar to the findings of [12, 24, 25] where the addition of more capsule layers has either led to stagnated or decreased performance. We strongly feel that our work should be revisited when a capsule network that can both scale and retain the desirable properties of capsule networks is found.
As stated in the original dataset proposal, the problem setting and methodology presented relies on the group elements being known. Hence, the applicability of the proposed method is only possible in settings where group elements are known. However, we present findings where alternative group elements are provided that suggests the CapsNets are more capable than previous invariant methods at capturing equivariant properties, thus opening an intriguing direction for future work.
Acknowledgments and Disclosure of Funding
We would like to thank the University of Aberdeen’s HPC facility for enabling this work.
@@
We use the 3DIEBench [14] dataset [^1] The full dataset and splits employed can be found at https://github.com/facebookresearch/SIE, which has been created specifically to be a hard yet controlled test-bed for invariant and equivariant methods. The dataset consists of 52,472 3d objects across 55 classes of 3d objects from ShapeNetCorev2 [6] posed in 50 different views as well as the latent information of the view, this can be seen in figure 7. For training we then randomly select two views from each model in the training set. The parameters by which the model could have been augmented can be seen in table 5.
Table 5: Values of the factors of variation used for the generation of 3DIEBench. Each value is sampled uniformly from the given interval. Object rotation is generated as Tayt-Bryan angles using extrinsic rotations. Light position is expressed in spherical coordinates. This table is sourced from [14].

### v3

@@ 4.1 Training Protocol
To directly compare with prior works employing the 3DIEBench dataset, we follow an identical training protocol, as defined in (Garrido et al., 2023). All methods employ a ResNet-18 encoder network ($f_{\theta}$), for the projection head ($h_{\phi}$) we compare a variety of hyperparameterisations, which we later describe in the following sections. For primary bench-marking we train our model for 2000 epochs using the Adam (Kingma & Ba, 2014) optimiser with default settings, a fixed learning rate of 1e-3 and batch size of 1024. For ablations and sensitivity analyses we train for 500 epochs and employ a batch size of 512, with other settings remaining unchanged. We find in practice that 500 epochs presents a strong correlation of performance. For all evaluations, pre-training was done with the equivariant criterion optimising for viewpoint rotation transformations, full details on these transformation groups, and the criterion are given in prior sections. By default the objective function weighting are as follows, $\lambda_{\text{inv}}=0.1,\;\lambda_{\text{equi}}=5,\;\lambda_{V}=10,\;\lambda_{C}=1$. Each self-supervised 2000 epoch pretraining run took approximately 22 hours using 3 Nvidia A100 80GB GPU’s, with 64 capsule models taking approximately 25 hours using 6 Nvidia A100 80GB GPU’s. All eval tasks are completed on a single Nvidia A100 80GB GPU and take approximately 6 hours for angle prediction, and 3 hours for classification.
Table 1: Evaluation of invariant properties via downstream classification task. Representations are learnt under the invariance and rotation equivariant objective, we evaluate both the representations and the intermediate embeddings of the projection head under varying number of capsules. FLOPs and Parameters correspond to computation during training, ‘-’ refers to non-compatible experiments. For non-capsule models, ‘All’, ‘Inv.’, and ‘Equi.’ refer to the full vector representation, the "left" half representation optimised under the invariant objective, and the "right" half representation optimised under the equivariant objective respectively. For capsule networks, ‘Inv.’, and ‘Equi.’ refer to the activation vector and pose matrices respectively.
@@
SIE 20.1M 13.07G 512 512 82.94 82.08 80.32
CapsIE - 16 12.7M 3.49G 16 256 76.51 - -
CapsIE - 32 14.7M 4.57G 32 512 79.14 - -
CapsIE - 64 20.4M 8.69G 64 1024 79.60 - -
@@
CapsIE - 16 12.7M 3.49G 16 256 - 74.90 -
CapsIE - 32 14.7M 4.57G 32 512 - 78.60 -
CapsIE - 64 20.4M 8.69G 64 1024 - 79.24 -
@@ 4.2 Downstream Evaluation
To evaluate the quality of representations learnt under the invariant and viewpoint rotation equivariant self-supervised criterion, we use the standard benchmark approach of learning downstream task specific networks with frozen representations as input. In our case we evaluate the representations in three distinct tasks to evaluate both invariant and equivariant properties, we use a linear evaluation training protocol of the frozen representations. Further details of the evaluation protocol are given in the appendix.B.2.
@@
Equivariant Evaluation. Evaluating equivariant properties is achieved through a rotation prediction task in which a three layer MLP is trained to predict the quaternions defining the rotation between two views of the same object. We train for 300 epochs using MSE loss. Similarly to rotation prediction, we evaluate the representation’s equivariant properties by regressing the the colour hue of an object view. We train a single linear layer for 50 epochs using MSE. For evaluating the performance of predicting equivariance, we use the metric $R^{2}=1-\frac{\sum_{i}(y_{i}-\tilde{y}_{i})^{2}}{\sum_{i}(y_{i}-\bar{y})^{2}}$ where $\{y_{i}\}$ are the ground truth, target values of the equivariance, $\bar{y}$ is the mean value of these targets, and $\{\tilde{y}_{i}\}$ are the predictions. Higher $R^{2}$ indicates the model has a better fit for predicting equivariant transformations.
Table 2: Evaluation of equivariant properties via downstream rotation prediction (left) and colour prediction (right) tasks. Representations are learnt under the invariance and rotation equivariant objective, we evaluate both the representations and the intermediate embeddings of the projection head under varying number of capsules. ‘-’ refers to non-compatible experiments. For non-capsule models, ‘All’, ‘Inv.’, and ‘Equi.’ refer to the full vector representation, the "left" half representation optimised under the invariant objective, and the "right" half representation optimised under the equivariant objective respectively. For capsule networks, ‘Inv.’, and ‘Equi.’ refer to the activation vector and pose matrices respectively.
@@
SIE 0.73 0.23 0.73 0.07 0.05 0.02
CapsIE - 16 0.68 - - 0.02 - -
CapsIE - 32 0.74 - - 0.04 - -
CapsIE - 64 0.72 - - 0.01 - -
@@
CapsIE - 16 - - 0.64 - - -0.01
CapsIE - 32 - - 0.71 - - -0.04
CapsIE - 64 - - 0.67 - - -0.09
Representation Quality. The performance of CapsIE for both invariant and equivariant benchmark tasks is given in Tables 1 and 2 respectively. We evaluate both the representations produced by the ResNet-18 encoder and the intermediate embeddings of the capsule layer projection head given different values for the numbers of capsules. We observe that across all models that the invariant properties captured within the representations marginally suffer compared to the MLP projector of SIE. This observation is expected given the significantly reduced number of embeddings employed in the invariant criterion compared to SIE. However, the evaluation of equivariant properties captured by the representations demonstrates that the use of a capsule projector in place of a MLP can lead to vastly improved performance in rotation prediction ($\uparrow 0.01\;R^{2}$) advancing the prior state-of-the-art whist also approaching the supervised baseline.
Intermediate Projector Embeddings. The role of the projector is primarily employed to decorrelate the embeddings on which the objective function operates on from the representations employed downstream. The premise is to avoid representations that are over-fit to the self-supervised objective (Bordes et al., 2022). However, it has been well studied that it can be beneficial to maintain a number of projector layers and instead utilise intermediate projector embeddings for downstream tasks. Specifically, the equivariant information that has shown to be captured in the object pose (Ribeiro et al., 2022).
We evaluate the intermediate embeddings output from the primary capsule layer in the same manner as the representations, however the activations and pose are given over a spatial region we perform average pooling to return an activation vector and a $4\times 4$ pose for each capsule which we then flatten into a vector. As with the representation evaluation we report our invariant and equivariant task performance in Tables 1 and 2 respectively. We find across all settings that evaluating the intermediate embeddings of the capsule projector leads to improved performance on rotation equivariant tasks compared to SIE intermediate embeddings. In our case, we demonstrate that the preservation of capsule layers for downstream tasks results in significantly closer evaluation performance to the representations, suggesting that capsule nets to presever such properties of interest compared to MLPs. Further investigations into this phenomena are left to future research.
@@ 4.3 Quantitative Evaluation of Equivariance
To quantitatively evaluate the equivariant performance of our method and capsule projector, we provide evaluations in line with those proposed in (Garrido et al., 2023). We report the Mean Reciprocal Rank (MRR) and Hit Rate at k (H@k) on the multi-object setting. Given a source and target pose of an object, we first compute the embeddings of each image, and pass the source embedding through the predictor and use the resulting vector to retrieve the nearest neighbours.
The Prediction Retrieval Error (PRE) gives an evaluation of predictor quality, and is given by the distance between its rotation $q_{1}\in\mathbb{H}$ and the target rotation $q_{2}$ as $d=1-<q_{1},q_{2}>^{2}$ of the nearest neighbour of the predicted embedding averaged over the whole dataset.
All the results evaluated by the aforementioned metrics are given in Table 3. Our CapsIE network outperforms EquiMod, Only Equivariance and SIE by a considerable margin across all metrics and for all dataset splits. We achieve strong perfromance on PRE, reporting 0.21 PRE on the validation set compared to 0.48 for EquiMod and Only Equivariance and 0.29 for SIE. The same large gains in equivariant performance are shown for MRR and H@1 and H@5. Note, a random H@1 results in a performance for 2% (0.02) demonstrating that our method lies well above random.
Table 3: Quantitative evaluation of the predictor when using a Capsule network projector, using PRE, MRR and H@k. The source dataset which embeddings are computed and the dataset used for retrieval are given in the format source-retrieval for PRE and source for MRR and H@k. Here source refers to the set from which embeddings are computed (train or val), while retrieval corresponds to the set used for comparison/retrieval (train, val, or all = train + val).
@@
CapsIE 0.17 0.21 0.20 0.60 0.47 0.50 0.36 0.71 0.58
@@ 4.4 Number of Capsules
Each capsule, in theory, should represent a unique concept, thus when the number of capsules is increased, logically so should the networks representation ability to capture an increasing number of semantic concepts. Observing the invariant performance during training (Figure 4) and downstream evaluation in Table 1, CapsIE gains slight improvement with the addition of more capsules. This demonstrates that our model has better utilised the additional representational power to improve performance. This pattern is also observed when evaluating equivariant properties (shown in Figure 4), yet is less pronounced. However, in Table 1 we also show that increasing the number of capsules of a supervised SR-Caps model trained in a standard supervised fashion is not an indicator of increased performance, aligning with prior capsule research (Everett et al., 2023). This differentiation in behaviour provides an interesting direction for future research.
Figure 4: Invariant and equivariant performance of encoder representations during training when varying number of capsules. (left) Classification evaluation performance (top-1 %) and (right) Rotation prediction online evaluation performance ($R^{2}$) both learned encoder representations. This evaluation is co-optimised during training and hence differs from the downstream evaluation reported in the Table 1 and 2. We observe that as capsule numbers increase so does performance, however at low capsule numbers later in training invariant objective takes precedence.
@@ 5.1 Equivariant Self-Supervised Learning
Self-supervised learning has seen the majority of its success in the invariant setting by either contrastive (Chen et al., 2020), information maximisation (Zbontar et al., 2021; Bardes et al., 2021), or clustering methods (Caron et al., 2021; Assran et al., 2022). All families of approaches rely on training a network to be invariant to transformations by increasing the similarity between embeddings of the same image under augmentation. The differing approaches emerge from alternative methods to avoid collapse, the phenomena where embeddings fall into a lower-dimensional subspace rather than the entire available embedding space resulting in a trivial solution (Hua et al., 2021). Although these methods differ, they all produce similarly performing representations, hence we employ information maximisation methods as the basis of this work due to their computational efficiency.
Learning to be invariant to transformations is typically useful for semantic discrimination tasks, yet preserving information about the transformations can be highly beneficial. Some approaches have attempted to capture specific information regarding transformations by predicting the applied augmentation parameters (Lee et al., 2021), preserving the strength of augmentations (Xie et al., 2022) and introducing rotational transformations (Dangovski et al., 2022). However, as stated in (Garrido et al., 2023), these methods provide no guarantee that a mapping is learnt in the latent space that reflects the transformations in the input space. Hence, methods have been employed that address this limitation, (Devillers & Lefort, 2022; Park et al., 2022; Garrido et al., 2023) all employ predictor networks to predict the displacement representations in the latent space given one view representation and the transformation parameters. The latter, SIE (Garrido et al., 2023), is the basis of our work, which further extends prior methods by splitting representation vectors into invariant and equivariant parts to better separate differing information.
@@ 5.2 Capsule Networks
CapsNets present an alternative architecture to CNNs, addressing their limitations by explicitly preserving hierarchical spatial relationships between features (Sabour et al., 2017). CapsNets replace scalar neurons with vector or matrix poses, representing specific concepts at different levels of a parse tree as the network goes deeper. The first layer (primary capsules) corresponds to the most basic parts, while capsules in deeper layers represent more complex concepts made up of the simpler concepts as they get closer to the final layer where each capsule corresponds to a specific class.
@@
The key novelty in CapsNets is the routing mechanism, which determines the contributions of lower-level capsules to higher-level capsules. Numerous routing algorithms, both iterative and non-iterative, have been proposed to address the efficiency and effectiveness of this process (Sabour et al., 2017; Mazzia et al., 2021; Feng et al., 2024; Ribeiro et al., 2020; Hinton et al., 2018; De Sousa Ribeiro et al., 2020; Yang et al., 2021; Everett et al., 2024). Among these, SRCaps (Hahn et al., 2019) introduces a non-iterative routing mechanism. This method maintains all of the desirable properties of CapsNets, such as equivariance while largely mitigating the time cost of iterative methods, at the price of a small amount of performance. However, SRCaps faces the same limitations as other CapsNets where high resolution or high class datasets are beyond the network’s abilities when trained in a standard fashion. For a more detailed description of capsule routing mechanisms please check this review (Ribeiro et al., 2022).
@@ 6 Conclusion
Our proposed method demonstrates how self-supervised CapsNets can be employed to better learn equivariant representations, leveraging architectural assumptions removing the the need to explicitly split representation vectors and train separate projector networks. The resulting solution, CapsIE, achieves state-of-the-art performance in equivariant downstream benchmarks with a improvement of 0.01 $R^{2}$ on prior self-supervised rotation prediction tasks. Our results contribute significantly to the avocation of CapsNets in self-supervised representation learning, introducing desirable properties with improved effectiveness over MLP projectors.
Ethics Statement. This work aims to learn higher quality and more applicable representations of images without human generated annotations, therefore such methods can lead to positive societal impacts the development of more accurate or informative models for a number of downstream tasks. However, as is the case with all vision systems, there is potential for exploitation and security concerns and one should take into consideration AI misuse when extending our method.
In addition, the use of CapsNets has shown little improvement when scaled, similar to the findings of (Everett et al., 2023; Mitterreiter et al., 2023; Nair et al., 2021) where the addition of more capsule layers has either led to stagnated or decreased performance. We strongly feel that our work should be revisited when a capsule network that can both scale and retain the desirable properties of capsule networks is found.
As stated in the original dataset proposal and the problem setting the methodology presented rely on the group elements being known. Hence, the applicability of the proposed method is only possible in settings where group elements are known. However, we present findings where additional group elements are preserved without prior knowledge of the group, this suggests the CapsNets are more capable than previous invariant methods at capturing equivariant properties, thus opening an intriguing direction for future work.
Reproducibility Statement. We have included the source code required to reproduce all experiments and ablation studies presented in the main body of the paper and the appendix. We also provide the details of our training protocol in the appendix.
@@
We use the 3DIEBench (Garrido et al., 2023) dataset [^1] The full dataset and splits employed can be found at https://github.com/facebookresearch/SIE, which has been created specifically to be a hard yet controlled test-bed for invariant and equivariant methods. The dataset consists of 52,472 3d objects across 55 classes of 3d objects from ShapeNetCorev2 (Chang et al., 2015) posed in 50 different views as well as the latent information of the view, this can be seen in figure 5. For training we then randomly select two views from each model in the training set. The parameters by which the model could have been augmented can be seen in table 4.
Table 4: Values of the factors of variation used for the generation of 3DIEBench. Each value is sampled uniformly from the given interval. Object rotation is generated as Tayt-Bryan angles using extrinsic rotations. Light position is expressed in spherical coordinates. This table is sourced from (Garrido et al., 2023).